OpenStudy (anonymous):

p4 - 1 factor completely

OpenStudy (anonymous):

p^4 - 1 =(p^2-1)(p^2+1) = (p+1)(p-1)(p^2+1)

OpenStudy (anonymous):

u r really helpin me out raheen thank u